Prove that if x is real, the expression (x-a)(x-c) / (x-b) is capable of assuming all values if a>b>c or a

Grade:12

1 Answers

Ramesh V
14 years ago

put Y = (x-a) (x-b) / (x-c)

so,its

x2 -(a+c+y)x + (ac+by) =0

now for roots to be real, its discriminant be positive

D>0 or (a+c+y)2 -4(ac+by) > 0

(a+c)2 + y2 +2y(a+c) -4ac -4by > 0

y2 + 2y(a+c-2b) + (a-c)2 > 0 ......(1)

here for any value of y(belongs to R) we have : y2  and  (a-c)2 always positive,

so for eqn (1) to be positive for all y(-ve or +ve) , 2(a+c-2b) shld be zero

i.e., a+c-2b = 0  or a,b,c in A.P series from where we can say that

either a<b<c  or c<b<a

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch